OpenStudy (nastech):

integral 1 divided by 1+sinx dx

OpenStudy (australopithecus):

Do you mean? \[\int\limits_{}^{}1dx/(1+\sin(x))\]

OpenStudy (nastech):

yes please

myininaya (myininaya):

\[\int\limits_{}^{}\frac{1}{1+\sin(x)} \cdot \frac{1-\sin(x)}{1-\sin(x)} dx=\int\limits_{}^{}\frac{1-\sin(x)}{\cos^2(x)} dx\] Did you try this?

OpenStudy (nastech):

no

OpenStudy (nastech):

caan you please continue with it for me. I'm an amateur with these things

myininaya (myininaya):

Break into two fractions

myininaya (myininaya):

write in terms of sec(x) and tan(x) and you will see it is not too terrible

OpenStudy (sburchette):

\[\int\limits(1-\sin x)dx/\cos ^2x\] Dividing (cosx)^2 into numerator we get, \[\int\limits (\sec ^2x-\tan x \sec x)dx \] Each term can be integrated independently. \[\tan x-\sec x+C\]